40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 21 15:13:44 2017, gautam, Update, IMC, IMC input beam mode matching IMG_7220.JPGIMC_ModeMatch.pdfsingleLensSensitivity.pdfsensitivity.pdfIMCmodeMatch.m
    Reply  Tue Mar 21 21:59:48 2017, gautam, Update, IMC, IMC input beam mode matching IMG_3654.JPGIMG_3655.JPG
       Reply  Wed Mar 22 00:33:00 2017, gautam, Update, IMC, IMC length offset nulling offsetInvestigation.pdfoffset_summing_amp.pdf
          Reply  Wed Mar 22 16:58:25 2017, gautam, Update, IMC, WFS sensing matrix measurements 
             Reply  Thu Mar 23 01:44:53 2017, gautam, Update, IMC, WFS sensing matrix measurements IMC_WFS_segment_TF.pdfIMC_WFS_channels_TF.pdfTFsummary.pdfIMC_WFS_170322.xlsx.zip
                Reply  Thu Mar 23 08:43:11 2017, rana, Update, IMC, WFS sensing matrix measurements 
                   Reply  Thu Mar 23 23:38:58 2017, gautam, Update, IMC, MC SUS damping gains stepped down 
                      Reply  Fri Mar 24 11:26:57 2017, gautam, Update, IMC, MC SUS damping gains restored 
                         Reply  Fri Mar 24 19:04:18 2017, gautam, Update, IMC, Seismic feedforward and WFS MCL_comparison.pdfseis_comparison.pdf
Message ID: 12904     Entry time: Fri Mar 24 11:26:57 2017     In reply to: 12903     Reply to this: 12906
Author: gautam 
Type: Update 
Category: IMC 
Subject: MC SUS damping gains restored 

I've restored the damping loop gains to their nominal values. Analysis of the coherence between MCL and seismometer channels under this reduced gain setting is underway, results to follow.

ELOG V3.1.3-